# MedHal: An Evaluation Dataset for Medical Hallucination Detection

### v1

@@ Abstract
We present MedHal, a novel large-scale dataset specifically designed to evaluate if models can detect hallucinations in medical texts. Current hallucination detection methods face significant limitations when applied to specialized domains like medicine, where they can have disastrous consequences. Existing medical datasets are either too small, containing only a few hundred samples, or focus on a single task like Question Answering or Natural Language Inference. MedHal addresses these gaps by: (1) incorporating diverse medical text sources and tasks; (2) providing a substantial volume of annotated samples suitable for training medical hallucination detection models; and (3) including explanations for factual inconsistencies to guide model learning. We demonstrate MedHal’s utility by training and evaluating a baseline medical hallucination detection model, showing improvements over general-purpose hallucination detection approaches. This resource enables more efficient evaluation of medical text generation systems while reducing reliance on costly expert review, potentially accelerating the development of medical AI research.
@@ 1 Introduction
Despite significant advancements in the field of Natural Language Processing, Large Language Models (LLMs) are prone to generating hallucinations (Huang et al. (2025)). This tendency poses a substantial obstacle to their application in multiple domains like medicine. While a robust hallucination detection is crucial, current hallucination metrics suffer from limitations in accuracy (Lin (2004); Banerjee & Lavie (2005); Zhang et al. (2019)). Traditional approaches relying on n-gram overlap (Lin (2004)) or entity overlap (Adams et al. (2024)) fail to capture the nuances of factual correctness. While more recent methods leverage general-purpose neural models (Kim et al. (2024); Laban et al. (2021)), incorporating context for potentially improved efficiency and have shown to correlate better with human judgment than n-gram overlap metrics (Zheng et al. (2023)), these models still exhibit suboptimal evaluation performance when applied to more specialized domain. Indeed, these models are mainly trained to detect general inconsistencies such as factual hallucinations or reasoning fallacies. In contexts like medical text generation, the most efficient way to detect hallucinated content is still to review generations by medical experts (Searle et al. (2023); Hegselmann et al. (2024)). However, this approach is costly and time-consuming, placing a significant burden on medical professionals. Therefore, a dedicated medical hallucination evaluation dataset is highly desirable. Such a resource could not only provide an accurate and relevant assessment of medical text generation, but also significantly reduce the cost and effort associated with medical AI research, facilitating its wider adoption. Current medical datasets often focus on single tasks like Question Answering (QA) or Natural Language Inference (NLI) to evaluate the hallucination levels of LLMs, limiting their scope and applicability to diverse medical text generation scenarios (Romanov & Shivade ; Pal et al. (2023); Chen et al. (2024); Yan et al. (2024); Agarwal et al. (2024); Hegselmann et al. (2024)). Furthermore, well-crafted medical hallucination datasets that currently exist only contain a few hundred samples, making them unsuitable for training large language models (LLMs) (Hegselmann et al. (2024)).
@@ 2 Related Work
Current state-of-the-art hallucination detection in the medical domain faces significant challenges due to the scarcity of comprehensive and diverse datasets. Existing hallucination datasets often focus on specific tasks or are limited in size. For instance, Med-Hallmark (Chen et al. (2024)) and Med-HVL (Yan et al. (2024)) focus on visual question-answering, assessing the ability of models to answer questions about medical images. MedHALU (Agarwal et al. (2024)) and Med-Halt (Pal et al. (2023)), on the other hand, are centered around QA and focus on evaluating LLMs’ ability to answer complex medical questions and detect inconsistencies in their responses. While the Hallucinations-MIMIC-DI dataset (Hegselmann et al. (2024)) provides a comprehensive annotation of discharge summaries and BHC sections, it only contains 100 samples, which is not sufficient to train a model or effectively evaluate another. Finally, some large datasets datasets like MedNLI exist for NLI tasks (Romanov & Shivade ). However, as shown in Figure 1, they rely on short, single-sentence premises and hypotheses, limiting their ability to capture complex medical reasoning and hallucinations.
@@ 3.1 Unified Task Formulation
To construct the dataset, a unified task formulation was established. Drawing inspiration from the NLI paradigm, each sample in the dataset is composed of a statement (analogous to the NLI hypothesis). This statement is classified as either factual or non-factual. The statement may be presented in the context of specific medical information (e.g., ”This patient has suffered from a myocardial infarction”) or as a general assertion within the medical domain (e.g., ”Aging causes an increase in blood pressure”). For each statement, the dataset provides a binary label indicating its factuality and, in the case of non-factual statements, an explanation detailing the specific inconsistency. A statement is defined as factual if all information contained within it is verifiable through the provided context or by established medical knowledge. Each sample in MedHal is structured as follows:
@@
Table 1: Example of samples are used to generate statements for each task
@@ 3.2 Question Answering Dataset Transformation
Question-answering (QA) datasets are structured around the presentation of a query followed by a set of potential responses, including binary (yes/no/maybe) and multiple-choice (A, B, C, …) formats. To generate factual samples from QA datasets, the question and its corresponding correct answer are transformed into a declarative statement using a large language model. Conversely, hallucinated samples are produced by pairing the question with incorrect answer options and subsequently converting these combinations into statements via the same LLM. Table 2 shows an example of a multiple-choice question converted into a statement. For generation, a consistent prompt template is employed for both factual and hallucinated sample generation. The prompt is given in Appendix 3. One-shot prompting is utilized to guide the LLM in accurately converting question-answer pairs into coherent statements. As for the explanation, certain QA datasets provide an explanation of why an answer is false. In these cases, we use the provided explanation for the non-factual statement. In other cases, we simply use the statement of the factual sample as the explanation.

### v2

@@ Abstract
We present MedHal [^1] MedHal is distributed under Apache 2.0 license. Link will be provided uppon acceptance., a novel large-scale dataset specifically designed to evaluate if models can detect hallucinations in medical texts. Current hallucination detection methods face significant limitations when applied to specialized domains like medicine, where they can have disastrous consequences. Existing medical datasets generally do not focus on hallucination detection and are either too small, containing only a few hundred samples, or focus on a single task like Question Answering or Natural Language Inference. MedHal addresses these gaps by: (1) incorporating diverse medical text sources and tasks; (2) providing a substantial volume of annotated samples suitable for training medical hallucination detection models; and (3) including explanations for factual inconsistencies to guide model learning. We demonstrate MedHal’s utility by training and evaluating a baseline medical hallucination detection model, showing improvements over general-purpose hallucination detection approaches. This resource enables more efficient evaluation of medical text generation systems while reducing reliance on costly expert review, potentially accelerating the development of medical AI research.
MedHal: An Evaluation Dataset for Medical Hallucination Detection
Gaya Mehenni1,2,3, Fabrice Lamarche1,2,3, Odette Rios-Ibacache 4,5, John Kildea4,5, Amal Zouaq1,2,3 1LAMA-WeST Lab, 2Polytechnique Montréal 3Mila - Quebec AI Institute 4Medical Physics Unit - RI-MUHC 5McGill University Correspondence: gaya.mehenni@polymtl.ca
@@ 1 Introduction
Despite significant advancements in the field of Natural Language Processing, Large Language Models (LLMs) are prone to generating hallucinations Huang et al. (2025). This tendency poses a substantial obstacle to their application in multiple domains like medicine. While robust hallucination detection is essential, existing hallucination metrics continue to suffer from accuracy limitations Lin (2004); Banerjee and Lavie (2005); Zhang et al. (2019). Traditional methods based on n-gram overlap Lin (2004) or entity overlap Adams et al. (2024) often fail to capture the nuances of factual correctness. Although more recent approaches employ general-purpose neural models Kim et al. (2024); Laban et al. (2021) and leverage contextual information for potentially improved evaluation, showing stronger correlation with human judgment compared to n-gram-based metrics Zheng et al. (2023), their performance remains suboptimal in specialized domains. These models are primarily trained to detect broad inconsistencies, such as factual inaccuracies or reasoning errors.
In high-stakes contexts like medical text generation, the most reliable method for identifying hallucinated content remains expert human review Searle et al. (2023); Hegselmann et al. (2024b). However, this process is both time-consuming and resource-intensive, placing a considerable burden on medical professionals. Consequently, there is a clear need for a dedicated medical hallucination evaluation dataset—both to assess the performance of LLMs on hallucination detection tasks and to enable fine-tuning for improved hallucination detection capabilities.
Currently, many medical datasets focus on narrow tasks such as Question Answering (QA) or Natural Language Inference (NLI), limiting their utility across the full range of medical text generation scenarios Romanov and Shivade ; Chen et al. (2024); Yan et al. (2024); Hegselmann et al. (2024b). Moreover, existing domain-specific hallucination datasets typically contain only a few hundred samples, rendering them insufficient for training or robustly evaluating large language models (LLMs) Hegselmann et al. (2024b). In this work, we propose MedHal, a dataset for the accurate and domain-relevant assessment of generated medical content, significantly reducing the cost and effort required for evaluation—ultimately facilitating broader adoption of medical AI systems.
@@ 2 Related Work
Current state-of-the-art hallucination detection in the medical domain faces significant challenges due to the scarcity of comprehensive and diverse datasets. Existing hallucination datasets often focus on specific tasks or are limited in size. For instance, Med-Hallmark Chen et al. (2024) and Med-HVL Yan et al. (2024) focus on visual question-answering, assessing the ability of models to answer questions about medical images. MedHaluAgarwal et al. (2024) and Med-Halt Pal et al. (2023), on the other hand, are centered around QA and focus on evaluating LLMs’ ability to answer complex medical questions and detect inconsistencies in their responses. While the Hallucinations-MIMIC-DI dataset Hegselmann et al. (2024b) provides a comprehensive annotation of discharge summaries and BHC sections, it only contains 100 samples, which is not sufficient to train a model or effectively evaluate another. Finally, some large datasets like MedNLI exist for NLI tasks Romanov and Shivade . However, as shown in Figure 1, they rely on short, single-sentence premises and hypotheses, limiting their ability to capture complex medical reasoning and hallucinations.
@@ 3.1 Unified Task Formulation
To construct the dataset, a unified task formulation was established. Drawing inspiration from the NLI paradigm, each sample in the dataset is composed of a statement (analogous to the NLI hypothesis). This statement is classified as either factual or non-factual. The statement may be presented in the context of specific medical information (e.g., "This patient has suffered from a myocardial infarction") or as a general assertion within the medical domain (e.g., "Aging causes an increase in blood pressure"). For each statement, the dataset provides a binary label indicating its factuality and, in the case of non-factual statements, an explanation detailing the specific inconsistency. A statement is defined as factual if all information contained within it is verifiable through the provided context or by established medical knowledge. Each sample in MedHal is structured as follows:
@@
Table 1: Example of samples used to generate statements for each task
@@ 3.2 Question Answering Dataset Transformation
Question-answering (QA) datasets are structured around a query followed by a set of potential responses, including binary (yes/no/maybe) and multiple-choice (A, B, C, …) formats. To generate factual samples from QA datasets, the question and its corresponding correct answer are transformed into a declarative statement using a large language model. Conversely, hallucinated samples are produced by pairing the question with incorrect answer options and subsequently converting these combinations into statements via the same LLM. Table 11 in Appendix A shows an example of a multiple-choice question converted into a statement. For generation, a consistent prompt template is employed for both factual and hallucinated sample generation. The prompt is given in Appendix 2. One-shot prompting is utilized to guide the LLM in accurately converting question-answer pairs into coherent statements. As for the explanation, certain QA datasets provide an explanation of why an answer is false. In these cases, we use the provided explanation for the non-factual statement. In other cases, we simply use the statement of the factual sample as the explanation.
3.3 Information Extraction Dataset Transformation
Information extraction (IE) datasets comprise, for each source document (e.g., clinical notes, clinical trials), a set of extracted text sequences related to specific concepts within that document. These extracted sequences, or "extractions," represent information pertaining to a particular attribute of the document. For example, from a clinical note, an extraction might represent the patient’s reason for the visit. To generate factual samples, the source document is used as the context, and the corresponding extraction is treated as the statement. Non-factual samples are generated by randomly interchanging extractions of the same value type between different documents (e.g., medications are swapped between patient records). When extractions are structured as key-value pairs, a large language model (LLM) is employed to transform these pairs into declarative statements. For instance, the key-value pair "visit motivation: Lower back pain" is converted into the statement "The patient’s visit motivation is lower back pain". This transformation is performed using the prompt template shown in the Appendix 3 within a one-shot prompting framework. The explanation for a non-factual statement is provided by the factual extraction associated with the same document and key concept. An example of how a sample from an information extraction dataset is transformed is shown in Table 12 in Appendix A. In cases where extractions are related to attributes with limited value diversity, such as sex, we ensure distinct values by explicitly switching them if random swapping results in identical values. The source documents are used without any preprocessing.
3.4 Natural Language Inference Dataset Transformation
3.5 Summarization Dataset Transformation
For summarization datasets, factual samples are constructed using genuine summaries or sentences extracted from them. Hallucinated samples are generated by extracting a single sentence from the original summary and employing an LLM to modify it, introducing self-contradictory information. The modified, hallucinated sentence is then reintegrated into the summary at its original position. The source text intended for summarization serves as the context, and the summary itself is treated as the statement. The explanation for a non-factual statement is simply the original sentence used to generate the hallucinated version. Sentences for hallucination generation are selected randomly, with a minimum length of 100 characters to ensure sufficient contextual information for the LLM. An example of how a sample is generated is shown in Table 13 in Appendix A. The prompt used to generate self-contradictory sentences is shown in Appendix 4. The goal of this process is to evaluate if models can effectively detect fine-grained hallucinations in statements consisting of long sequences of text.
Dataset Task S Content Type Source # Samples # Gen
MedMCQA QA Medical Content Pal et al. (2022) 183,000 278,021
MedNLI NLI Clinical Notes Herlihy and Rudinger (2021) 11,232 7,488
ACM IE Clinical Notes Bonnet and Boulenger (2024) 22,000 327,568
MedQA QA Medical Content Jin et al. (2020) 12,723 18,680
PubMedSum Sum Clinical Trials Gupta et al. (2021) 33,772 195,339
Table 2: Description of datasets, after filtering, used to create the MedHal benchmark (S is for Synthetic)
3.6 Model
For the creation of the MedHal dataset, we utilize the Llama-3-70B model Grattafiori et al. (2024), a state-of-the-art language model known for its strong performance across various language generation tasks. For all prompting strategies employed in the sample generation process, a one-shot learning approach was adopted to provide the model with examples of the desired output format prior to generation. See Sections 3.2, 3.3, 3.4, and 3.5 for details on what is generated for each task. Importantly, to ensure a fair evaluation of our proposed benchmark, the dataset creation was performed exclusively using the training splits of the individual source datasets. This approach allows for the independent evaluation of our trained models on the original test sets of the constituent datasets.
3.7 Data Quality Assessment
To ensure data quality, we incorporated a filtering step where a Llama-3.3-70B model assessed the factuality of each generated statement, given its context. For tasks with subtle nuances, such as QA and summarization, the model was provided with the correct answer or the original sentence, respectively, to guide its judgment. For NLI tasks, it simply evaluated the statement’s factuality. Samples were retained if the model’s assessment matched the dataset’s assigned label; otherwise, they were removed.
Table 3 presents the percentage of samples retained for each dataset after this filtering process. A majority of samples were retained across most datasets, with the exception of ACM, which is the only synthetic dataset in our collection.
Dataset % Samples Kept
SumPubMed 99.30
ACM 82.00
MedMCQA 96.87
MedQA 99.44
MedNLI 100.00
Table 3: Percentage of samples kept after the filtering step for each dataset.
3.8 MedHal Dataset Description
The MedHal benchmark is distributed in several configurations to accommodate diverse research needs. The full dataset comprises a total of 827,096 samples. Additionally, we provide a length-filtered (LF) version, retaining only samples where the combined length of the context and statement is less than 30,000 characters. This specific length constraint is chosen to align with the 8,192 token context length common to most current state-of-the-art 1-10B parameter models, reflecting our study’s focus on the performance of smaller language models in medical settings.
Furthermore, we release a balanced version (BAL) of the dataset in terms of the tasks. In this configuration, the Question Answering (QA), Summarization, and Information Extraction (IE) tasks each contain approximately 150,000 samples. The NLI task, however, could not be balanced to this extent due to the inherently limited number of available samples for that specific task.
Table 4 details the number of samples and their respective proportions for each task across the different dataset versions.
Task Full LF LF + BAL
Summarization 0.24 0.19 0.31
Information Extraction 0.40 0.42 0.31
Question Answering 0.36 0.38 0.35
NLI 0.01 0.01 0.02
Total samples 827,096 779,376 469,025
Table 4: Sample proportions for each task across different dataset versions (LF=Length Filtered, BAL=Balanced)
3.9 Expert Validation
To validate the quality of the filtered dataset, one of the authors (PhD student in Medical Physics), conducted a manual annotation of a subset of the dataset. We randomly selected 50 examples from each Task (25 for each Factual and Non-Factual labels [^2] We only select Non-Factual samples for the Summarization task since Factual Samples contain no modification of the data.). The annotator is then tasked with validating the label, considering both the Context and Statement. A sample is considered valid if its label is judged as correct by the annotator. Annotation guidelines are presented in Appendix D. Statistics describing the validity of each task is presented in Table 5. We observe near perfect validity for all tasks, and manual observation of the invalid examples show samples that lack relevance or important pieces of information from the context.
Task Valid samples (%)
Question Answering 0.91
Information Extraction 0.96
Natural Language Inference 0.98
Summarization 1.0
Table 5: Expert validation results per task.
Type Model Factuality Explanation
P R F1 BLEU R1 R2 N
General Llama-3.2-1B 0.51 0.41 0.46 0.01 0.08 0.03 1040
Llama-3-8B 0.66 0.58 0.62 0.03 0.12 0.08 15962
Medical BioMistral-7B 0.64 0.51 0.57 0.09 0.25 0.19 13365
MedLlama-8B 0.94 0.26 0.40 0.09 0.32 0.22 22395
OpenBioLLM-8B 0.61 0.76 0.68 0.01 0.04 0.01 7153
Evaluator Prometheus-2-8x7B 0.85 0.69 0.76 - - - -
HallOumi-8B 0.67 0.76 0.72 - - - -
Table 6: Performance of models on MedHal’s test set (N refers to the number of samples where an explanation was extracted when prompting the model)
4 Experiments
4.1 General Evaluation
To determine how well current AI models detect medical hallucinations, we evaluated their performance on the MedHal test set. This assessment helps us identify which models are most adept at recognizing hallucinated content within a medical context and, more importantly, provides insights into whether specific fine-tuning strategies improve medical hallucination detection. We compare the performance of general-purpose models (Llama-3.2-1B, Llama-3-8B), medical models fine-tuned on general medical datasets (BioMistral-7B, MedLlama-8B, OpenBioLLM-8B), and evaluator models fine-tuned to judge answers of other LLMs (Proemtheus-2-8x7B, HallOumi-8B) . For general-purpose and medical models, we used the prompt template detailed in Figure 5. For Prometheus-2-8x7B and HallOumi-8B, we adhered to the prompt formats recommended by their original authors, as these are optimized for their fine-tuned performance.
We use two main types of metrics to evaluate the models: factuality metrics and explanation metrics. Factuality metrics measure how accurately a model identifies factual versus non-factual content. This includes common measures like precision, recall, and F1-score. Due to inconsistencies in model output when using the prompt format from Figure 5. You can see these results in Table 6. Explanation metrics assess the validity of the explanations that the models provide for non-factual statements. Specifically, these metrics check if a model, after identifying non-factual content, correctly pinpoints the exact erroneous part of the statement. The explanation metrics are ROUGE-1 (R1), ROUGE-2 (R2) Lin (2004), and BLEU Papineni et al. (2002) scores. To ensure a valid comparison, we only consider samples where both the model’s prediction and the ground truth label indicate a non-factual statement. This guarantees that a true explanation exists and that the model attempted to generate one.
4.2 Impact of Specialized Training
We fine-tuned several models on the MedHal training set to assess the impact of specialized training on their performance. Our main goal was to see if fine-tuning a medical model specifically for medical hallucination detection, or specializing a general hallucination detection model on medical data, would yield better results. We also investigated whether simply fine-tuning a general-purpose model could achieve similar performance, potentially bypassing the need for more specialized initial training. For this experiment, we fine-tuned Llama-3-8B, Llama-3-OpenBioLLM-8B, and HallOumi-8B on MedHal. The prompt used during training is shown in Figure 6. All models were fine-tuned using the same QLora configuration [^3] For details on training setup, see Appendix C. We report the same metrics detailed in Table 6, along with the difference in F1 score between the fine-tuned and non-fine-tuned versions of each model in Table 7.
[…]
P R F1 BLEU R1 R2
Llama-3-8B 0.93 0.84 0.88 0.52 0.81 0.74 +0.26
OpenBioLLM-8B 0.90 0.88 0.90 0.53 0.81 0.74 +0.22
HallOumi-8B 0.92 0.80 0.86 0.51 0.80 0.72 +0.14
[…]
4.3 Task Ablation
We perform an ablation on the task training data to better understand the individual contribution of each task to the models’ overall performance. In this setting, we fine-tune the Llama-3-8B base model separately on the training samples for each task: Summarization, Question Answering, and Information Extraction. The same prompt format is used across all experiments to ensure consistency. We exclude the Natural Language Inference (NLI) task from this analysis because it contains significantly fewer samples compared to the others, which would make a direct comparison of its impact inequitable. By isolating each task, this ablation allows us to assess its relative importance. A substantial performance gain from a single task would indicate its high value, whereas a decrease or negligible improvement compared to the baseline could suggest the task is less critical for the overall goal. Results are shown in Table 8.
[…]
P R F1 BLEU R1 R2
Summarization 0.71 0.95 0.81 0.46 0.69 0.62 +0.19
Question Answering 0.78 0.67 0.72 0.19 0.50 0.38 +0.10
Information Extraction 0.73 0.72 0.74 0.17 0.54 0.41 +0.12
[…]
4.4 Downstream Performance
We assess the validity of our MedHal dataset and its utility for benchmarking by evaluating the downstream task performance of models fine-tuned on MedHal on MedNLI Herlihy and Rudinger (2021) and another hallucination detection dataset Hegselmann et al. (2024a).
4.4.1 MedNLI
We first evaluate base model on MedNLI and compare them against a Llama-3-8B model fine-tuned on MedHal. Additionally, we establish a baseline by fine-tuning a Llama-3-8B model on the MedNLI training set for 5 epochs to validate the impact of our tasks on performance. As models fine-tuned on our dataset cannot detect neutral statements in MedNLI, we filter out these samples from the test set during evaluation of all models. We report the F1 scores of all models in Table 9.
Fine-Tuning Base Model F1-Score
None Llama-3.2-1B 0.64
Llama-3-8B 0.65
BioMistral-7B 0.56
MedLlama-3-8B 0.66
OpenBioLLM-8B 0.64
Prometheus-2-8x7B 0.62
HallOumi-8B 0.89
MedNLI Llama-3-8B 0.95
MedHal Llama-3-8B 0.96
Table 9: F1-Score of models on the MedNLI test set (samples with a neutral labels have been filtered out)
4.4.2 Hallucination Dataset
This benchmark Hegselmann et al. (2024a) is derived from MIMIC-III and contains summaries of patient notes generated by frontier models, which were then annotated for factual inconsistencies by medical students. We consider a summary as non-factual if it contains any of the eleven types of factual inconsistencies in the dataset, including ’Fact Contradicted’, ’Unsupported Medication’, or ’Unsupported Procedure’. Although the dataset contains only 210 samples—152 of which include hallucinations—its recent release makes it highly unlikely that the evaluated models were exposed to it during training. We compare 1-shot prompting strategy using several models - as models consistently failed to identify hallucinations in a zero-shot setting-, a base Llama-3-8B model fine-tuned on the MedHal train set to measure the impact of our dataset, and evaluation models (Prometheus-2-8x7B, HallOumi-8B). Results are shown in Table 10.
Model Precision Recall F1
Llama-3.2-1B (1 shot) 0.81 0.32 0.45
Llama-3.1-8B (1 shot) 1.00 0.13 0.22
OpenBioLLM-8B (1 shot) 0.85 0.60 0.70
Prometheus-2-8x7B 0.77 0.75 0.76
HallOumi-8B 0.91 0.61 0.73
MedHal (Llama-3-8B) 0.95 0.63 0.76
Table 10: Evaluation of different models on an Hallucination Dataset
5 Discussion
5.1 Performance of Models
Our evaluation reveals several key insights into the performance of various models in detecting and explaining medical hallucinations. A notable finding is the clear superiority of models specifically fine-tuned as evaluators in the task of factuality detection. Prometheus-2-8x7B emerges as the most accurate model, achieving the highest F1-score of 0.76, followed closely by HallOumi-8B with an F1-score of 0.72. Surprisingly, general-purpose models demonstrate a strong aptitude for this task, with Llama-3-8B (F1=0.62) outperforming some specialized medical models. The performance of medical models is mixed; while OpenBioLLM-8B surpasses the general-purpose baseline with an F1-score of 0.68, other models in this category lag. When assessing the quality of explanations for non-factual content, however, the trend reverses. Here, medical models show a distinct advantage, with MedLlama-8B significantly outperforming all other models across explanation metrics (R1=0.32, R2=0.22) while having the highest number of valid explanations (indicating a strong capacity to follow instructions). We perform a detailed analysis of model errors in Appendix E.
5.2 Task-Specific Performance Variations
Table 8 shows that each task in the dataset contributes uniquely to improving model performance in detecting factual and non-factual content, with all tasks boosting the baseline F1 score by at least 10%. Among them, the Summarization task shows the most significant impact. Fine-tuning solely on Summarization leads to substantially greater gains over the baseline than fine-tuning on Question Answering or Information Extraction. These improvements are consistent across both "Factuality" and "Explanation" metrics. The effectiveness of Summarization can be attributed to its inherent complexity. Unlike Question Answering, which typically draws on general medical knowledge, Summarization requires a more complex understanding of the context and the statement. While Information Extraction also involves contextual analysis, it deals with short, focused statements. Summarization, in contrast, involves longer statements and minor inconsistencies, requiring the model to analyze both the statement and the broader context. This might lead to the model learning more robust, generalizable representations for hallucination detection, which also benefit tasks like Information Extraction and some Question Answering cases. Representations learned from Question Answering and Information Extraction do not transfer as effectively to other tasks in the dataset.
5.3 Evaluation on Downstream Datasets
The results on the MedNLI dataset, shown in Table 9, highlight the benefits of MedHal fine-tuning. Our Llama-3-8B model achieves an F1-score of 0.96, a 7% improvement over the best baseline (HallOumi-8B), and slightly outperforms the Llama-3-8B model fine-tuned solely on MedNLI (F1=0.95), despite only encountering MedNLI samples once during multi-task training, compared to 5 epoch for Llama-3-8B. This small margin may be due to performance saturation on MedNLI for models of this size, but it still underscores the value of diverse task signals. On the MIMIC-based hallucination dataset, our MedHal-tuned model achieves an F1 score of 0.76, matching Prometheus-2-8x7B despite being nearly six times smaller. While Prometheus balances precision (0.77) and recall (0.75), our model shows a high-precision profile, achieving 0.95 precision and 0.63 recall. In clinical settings, such high precision is highly valuable, ensuring flagged statements are rarely false positives. In contrast, models like Llama-3.1-8B (1-shot) achieve perfect precision but suffer from extremely low recall (0.13), limiting practical use. Our model also surpasses others like OpenBioLLM-3-8B (F1=0.70) and HallOumi-8B (F1=0.73), reaffirming the strength of MedHal’s multi-task training approach.
Conclusion
In this study, we introduced MedHal, the first large-scale dataset designed to specifically address the challenge of hallucination detection across diverse clinical tasks, including summarization, natural language inference, information extraction, and question answering. We evaluated a range of medical and general-purpose models on this dataset, providing a comprehensive benchmark for assessing their performance in this critical domain. Furthermore, we fine-tuned multiple models on MedHal to evaluate the impact of specialized pre-training on medical hallucination detection performance. Plus, we also assess which tasks models tend to struggle more with. MedHal represents a significant step forward in addressing the limitations of existing hallucination detection methods in the medical field.
Limitations
It is important to acknowledge some limitations of our methodology, the main one being that we rely on generated content. Ideally, a more comprehensive evaluation involving medical professionals could be conducted to validate the accuracy and clinical relevance of the generated dataset. We also note that while our expert validation showed great data quality, a multi-annotator setup would allow measurement of agreement metrics and would improve our validity claims. Plus, we could incorporate more complex multi-hop reasoning hallucination detection tasks into the dataset. This would allow us to identify error patterns and find more room for improvement in developing hallucination detection models. Finally, this resource is only available in English and augmenting it with other languages would be beneficial. This is left for future work.
Appendix
Appendix A Data Transformation Examples